The following are several transactions of Ardery Company that occurred during the current year and were recorded in permanent (that is, balance sheet) accounts unless indicated otherwise:

Date

Transaction

Apr. 1 Purchased a delivery van for $20,000, paying $1,000 down, and issuing a 1-year, 6% note payable for the $19,000 balance. It is estimated that the van has a 4-year life and an $800 residual value; the company uses straight-line depreciation. The interest on the note will be paid on the maturity date.
May 15 Purchased $850 of office supplies.
June 2 Purchased a 2-year comprehensive insurance policy for $720.
Aug. 1 Received 6 months' rent in advance at $250 per month and recorded the $1,500 receipt as Rent Revenue.
Sept. 15 Advanced $600 to sales personnel to cover their future travel costs.
Nov. 1 Accepted a $4,000, 6-month, 12% (annual rate) note receivable from a customer, the interest to be collected when the note is collected.

The following information also is available:

1. On January 1, the Office Supplies account had a $250 balance. On December 31, an inventory count showed $200 of office supplies on hand.
2. The weekly (5-day) payroll of Ardery Company amounts to $2,000. All employees are paid at the close of business each Wednesday. A 2-day accrual is required for the current year.
3. Sales personnel travel cost reports indicate that $500 of advances had been used to pay travel expenses.
4. The income tax rate is 30% on current income and is payable in the first quarter of next year. The pretax income before the adjusting entries is $8,655.

Required:

On the basis of the above information, prepare journal entries to record whatever adjustments are necessary to bring the accounts up to date on December 31.
Answer #1

Adjustment entries on 31st December

Date

Particulars

Debit

Credit

December 31

Interest expenses

          Outstanding interest expenses*

(provided for outstanding interest on $ 19,000 notes payable @ 6% p.a. for 9 months)

855

855

December 31

Depreciation expenses

            Accumulated depreciation

(provided depreciation for 9 months on delivery van)

3,600

3,600

December 31

Office supplies expenses

         Office supplies

(office supplies consumed)

900

900

December 31

Insurance expenses

       Prepaid insurance

(provided for 5 months insurance expired out of prepaid insurance)

210

210

December 31

Rent revenue

     Unearned rent revenue

(Provided for one month rent received in advance )

250

250

December 31

Travel expenses

     Prepaid travel expenses

(Provided for expired travel cost out of prepaid travel expenses)

500

500

December 31

Payroll expenses

     Payroll expenses payable

(Provided for outstanding payroll expenses for 2 days)

800

                        

800

December 31

Tax expenses

     Tax payable

(provided for tax payable @ 30% on pre-tax income of $ 8,655)

2,596.50

2,596.50

December 31

Accrued interest**

     Interest income

(provided for accrued interest on notes receivable of $ 4,000 @ 12% p.a. for 2 months)

80

80

* An alternative for outstanding interest expenses account can be notes payable account.

** An alternative for accrued interest can be notes receivable account.

Working notes

1. Depreciation on delivery van has been provided for 9 months which is as follows

Annual depreciation = (20,000 – 800)/4 = 4,800

Depreciation for 9 months = (4,800/12) * 9 = 3,600

2. Office supplies consumed = opening supplies + purchase of supplies – closing supplies

i.e. = 250 + 850 – 200 = 900

3. Insurance expenses has been calculated as follows

Out of 24 months insurance payment only 7 months insurance has expired, it has been provided as insurance expenses out of prepaid insurance expenses

Insurance expenses = (720/24 ) * 5 = $ 210

4. Out of six months rent received, five months rent has expired, it has been calculated as follows

Ret received expired = 250 * 5 = 1,250

Remaining 250 ( 1,500 – 1,250) has been recorded as rent revenue received in advance .

5. Payroll expenses for 2 days have been calculated as follows

Outstanding payroll expenses = (2,000/5) * 2 = 800

These have been provided as outstanding expenses